OpenStudy (anonymous):

simplify the expression. -6+i\-5+i

OpenStudy (jdoe0001):

use the conjugate of the denominator, to "simplify" or "rationalize" the denominator, thus getting rid of that pesky "i" so the denominator is -5+i it's conjugate is the same thing but with a negative in between, that is -5 - i so \(\bf \cfrac{ -6+i }{ -5+i }\times \cfrac{-5-i}{-5-i}\implies \cfrac{(-6+i)(-5-i)}{(-5+i)(-5-i)}\\ \quad \\ \textit{keeping in mind that }\qquad (a-b)(a+b) = a^2-b^2\qquad then\\ \quad \\ \cfrac{(-6+i)(-5-i)}{(-5+i)(-5-i)}\implies \cfrac{(-6+i)(-5-i)}{-5^2-i^2} \) multiply and add like-terms, see what you get